OpenStudy (anonymous):

homogeneous equations (x^2+y^2)dx+2xydy=0

OpenStudy (anonymous):

im using y=vx dy=xdv+vdx \[(x^2+(vx)^2)dx+2x(vx)(xdv+vdx)=0\] then \[x^2dx+v^2x^2dx+2vx^2(xdv+vdx)=0\] finally \[x^2dx+v^2x^2dx+2vx^3dv+2v^2x^2dx=0\] stuck here and not sure if this is even right....

OpenStudy (anonymous):

@Nurali does the work I did look formatted to you? normally it looks formatted for me but its not. Do you know how I can change it back?

OpenStudy (anonymous):

@Nurali thanks a lot!

OpenStudy (nurali):

Anytime.

OpenStudy (nurali):

Please click best response.